# StyleGuard: Preventing Text-to-Image-Model-based Style Mimicry Attacks by Style Perturbations

### v1

@@ A.9 Evaluation over Online Black-box Upscaler
We use an online black-box upscaler, Finegrain Image Enhancer (FIE) HuggingFace [2024], to further evaluate our method. For FIE, we use the default setting, with the upscale factor set as 2 and the ControlNet scale set as 0.6. The Gaussian noise standard deviation is set as 0.1. Figure 1 shows the results of fine-tuning using a Van Gogh-style image. The first row is the result of fine-tuning using a clean image. The second row is the result of fine-tuning using a protected image after purification and then using Dreambooth. When using FIE for purification, the style of the image is somewhat similar to Van Gogh’s style, but it is still not as good as the fine-tuning result of the clean image. We believe that this is because FIE has a high degree of denoising, which can purify noise to a certain extent, but will also modify the content of the image to a certain extent, such as the details of the brushstrokes, changes in lines, etc., thus causing the style of the fine-tuned image to change.
@@ B.1 Compare the Images Trained on Clean Images and Protected Images
Figure 10 and Figure 11 compare the results of style mimicry on clean and protected images with StyleGuard protection. For the StyleGuard, we generate perturbations using SD1.4 and SD x4 upscaler. During the test, we first apply the Noise Upscale using the SD x2 upscaler and then train the SD1.5 model on the protected paintings. With the protection, the protected image quality significantly decreases, and the style is also changed from the original images.
Figure 10: The results of style mimicry on clean images that are without any protection. We train the SD v1.5 model on Van Gogh’s paintings.
Figure 11: The results of style mimicry on protected images with the StyleGuard protection. For the StyleGuard, we generate perturbations using SD1.4 and SD x4 upscaler. During the test, we first apply the Noise Upscale using the SD x2 upscaler and then train the SD1.5 model on the protected paintings.

### v2

@@ A.9 Evaluation over Online Black-box Upscaler
We use an online black-box upscaler, Finegrain Image Enhancer (FIE) [1], to further evaluate our method. For FIE, we use the default setting, with the upscale factor set as 2 and the ControlNet scale set as 0.6. The Gaussian noise standard deviation is set as 0.1. Figure 1 shows the results of fine-tuning using a Van Gogh-style image. The first row is the result of fine-tuning using a clean image. The second row is the result of fine-tuning using a protected image after purification and then using Dreambooth. When using FIE for purification, the style of the image is somewhat similar to Van Gogh’s style, but it is still not as good as the fine-tuning result of the clean image. We believe that this is because FIE has a high degree of denoising, which can purify noise to a certain extent, but will also modify the content of the image to a certain extent, such as the details of the brushstrokes, changes in lines, etc., thus causing the style of the fine-tuned image to change.
@@ B.1 Compare the Images Trained on Clean Images and Protected Images
Figures 10 and Figure 10 compare the results of style mimicry on clean and protected images with the StyleGuard protection. For StyleGuard, we generate perturbations using SD1.4 and SD x4 upscaler. During the test, we first apply the Noise Upscale using the SD x2 upscaler and then train the SD1.5 model on the protected paintings. With protection, the quality of the protected image decreases significantly and the style is also changed from the original images.
Figure 10: The results of style mimicry on clean images without any protection. We train the SD v1.5 model on Van Gogh’s paintings.
Figure 11: The results of style mimicry on protected images that with the StyleGuard protection. For the StyleGuard, we generate perturbations using SD1.4 and SD x4 upscaler. During the test, we first apply the Noise Upscale using the SD x2 upscaler and then train the SD1.5 model on the protected paintings.